a collection of 81 coins consists of nickels and dinmes only and has a value of $5.80 find the number of nickels and dimes in the collection
Answer by checkley77(12844)

D+N=81 OR D=81-N
.10D+.05N=5.80
.10(81-N)+.05N=5.80
8.10-.10N+.05N=5.80
-.05N=5.80-8.10
-.05N=-2.30
N=-2.30/-.05
N=46 NICKELS.
D=81-46=35 DIMES.
PROOF:
.10*35+.05*46=5.80
3.50+2.30=5.80
5.80=5.80